# Disentangled Dynamic Intrusion Detection

### v2

@@ Abstract
Network-based intrusion detection system (NIDS) monitors network traffic for malicious activities, forming the frontline defense against increasing attacks over information infrastructures. Although promising, our quantitative analysis shows that existing methods perform inconsistently in declaring various known or unknown attacks (e.g., $31$% F1 for the Backdoor and $93$% F1 for DDoS by a GCN-based state-of-the-art method), and perform poorly in few-shot intrusion detections (e.g., dramatically drops from $91$% to $36$% in 3D-IDS, and drops from $89$% to $20$% in E-GraphSAGE). We reveal that the underlying cause is entangled distributions of flow features. This motivates us to propose DIDS-MFL, a disentangled intrusion detection method to handle various intrusion detection scenarios. DIDS-MFL involves two key components to address the performance inconsistency and few-shot detection performance deterioration, respectively: a double Disentanglement-based Intrusion Detection System (DIDS) and a plug-and-play Multi-scale Few-shot Learning-based (MFL) intrusion detection module. Specifically, the proposed DIDS first disentangles traffic features by a non-parameterized optimization based on mutual information, automatically differentiating tens and hundreds of complex features of various attacks. Such differentiated features will be fed into a memory model to generate representations, which are further disentangled to highlight the attack-specific features. Our DIDS additionally uses a novel graph diffusion method that dynamically fuses the network topology for spatial-temporal aggregation in evolving data streams. Furthermore, the proposed MFL involves an alternating optimization framework to address the entangled representations in few-shot traffic threats with rigorous derivation. MFL first captures multi-scale information in latent space to distinguish attack-specific information and then optimizes the disentanglement term to highlight the attack-specific information. Finally, MFL fuses and alternately solves them in an end-to-end way. To the best of our knowledge, DIDS-MFL takes the first step toward dynamic intrusion detection in practice under various scenarios. Equipped with DIDS-MFL, administrators can effectively identify various attacks in encrypted traffic, including known, unknown ones, and the few-shot threats that are not easily detected. Experiments show the superiority of our proposed DIDS-MFL. For few-shot NIDS, our DIDS-MFL achieves a $73.59$% - $139.53$% improvement in average F1-score over $11$ baselines and shows versatility in multiple baselines and multiple tasks. Our code is available at https://github.com/qcydm/DIDS-MFL
Index Terms:
Intrusion Detection, Network Security.
I Introduction
Unauthorized attempts like password cracking [1], man-in-the-middle attacks (MITM) [2], and denial-of-service (DoS) [3] are known as the network attacks[4], targeting an organization’s digital assets with the intention of data leakage or the execution of harmful deeds. These attacks are frequently perceived as network anomalies [5] due to their distinct characteristics that differ from standard traffic patterns. An alarming statistic [6] notes that $31\%$ companies across the globe experience a daily average of at least one cyber attack, a frequency amplified by the proliferation of mobile online business endeavors. The situations necessitate an intelligent system deployment to assist network administrators in the automated segregation of these anomalies from the vast sea of internet traffic. A network-based intrusion detection system (NIDS) [7], which monitors network traffic and identifies malicious activities, facilitates administrators to form the frontline defense against increasing attacks over information infrastructures (e.g., sensors and servers). Hence, NIDS is widely applied in many information systems of governments and e-commercial business sectors [8]. Figure 1 demonstrates how NIDS builds a frontline defense against two network attacks, protecting systems from potential cyber threats.
@@ I Introduction
Figure 1: Illustration of two network attacks DoS and MITM. DoS floods the target with massive traffic to overwhelm an online service, and MITM eavesdrops on the communication between two targets and steals private information. An NIDS can be easily deployed in a single location to collect statistical features and alert administrators for potential threats.
Existing NIDS can be categorized into two types, i.e., signature-based ones [9, 10, 11] and anomaly-based ones [12, 13, 14]. The former detects network attacks based on pre-defined patterns or known malicious sequences stored in a database, such as the number of bytes in traffic. These patterns in the NIDS are referred to as signatures. The latter anomaly-based NIDS learns to track attacks with machine learning techniques. Early statistical approaches [2, 3], such as Support Vector Machine (SVM), Logistic Regression (LR), and Decision Tree (DT), rely on carefully designed handcrafted features to learn classification boundaries. Recent deep learning-based methods [15, 16] use millions of neural parameters to mine the knowledge underlying the training samples, and have achieved great success in automatically modeling complex correlations for tens and thousands of features. The state-of-the-art E-GraphSAGE [17] employs graph convolution networks (GCNs) to learn the feature representations for better prediction.
@@
Figure 2: Quantitative analysis on CIC-TON-IOT. (a) Comparisons of detecting various attacks, which are regarded as an unknown type in evaluation. Specifically, we train an SVM model without using the data points of these attacks, and evaluate the instances of these attacks on the test set. (b) and (c) show the feature distributions of two attacks, MITM and DDoS, respectively. (d) Comparisons of detecting various known attacks on the previous state-of-the-art deep learning model E-GraphSAGE, (e) and (f) are correlation maps of representations of the two attacks, where the representations are generated by E-GraphSAGE.
Figure 3: (a) Performance comparisons among different intrusion detection methods with normal-size traffic and few-shot traffic. (b) F1-score comparisons between DIDS few-shot learning and two SOTA few-shot learning baselines. We repeat the above comparisons 10 times.
Although promising, existing NIDS approaches face two challenges in detecting traffic threats accurately:
@@
  • Challenge $1$. The existing NIDS methods yield extremely inconsistent results in identifying distinct attacks. For statistical methods, Figure 2 (a) demonstrates that the detection performance of an SVM-based method [18] for an unknown attack  [^1] The unknown attacks referred to in this paper are never appeared new type of attacks in training set can be as low as $9\%$ in terms of F1 on the CIC-ToN-IOT [19] dataset. While the model achieves $40\%$ F1 in declaring another unknown threat (DDoS) on the same benchmark. Regarding the deep learning-based methods, Figure 2 (d) demonstrates that E-GraphSAGE achieves a lower than $20$% F1 score for MITM attacks and a higher than $90$% F1 score for DDoS on the CTC-ToN-IOT dataset.
  • Challenge $2$. The existing NIDS approaches struggle to detect few-shot traffic threats accurately. In practical scenarios, the strengthening of defensive measures and the consequent escalating trend of cyber threats have inevitably caused the rise of new-type attacks with a limited instance number, referred to as few-shot traffic threats. However, our empirical observations reveal that the NIDS methods have dramatic performance deterioration in few-shot scenarios, as shown in Fig. 3 (a), e.g., our previous work DIDS [20] dramatically drops from $91.57$% under supervised learning setting to $36.12$% under the few-shot learning setting over the CIC-ToN-IoT dataset, and similarly, EURLER and E-GraphSAGE have an average F1-scores lower to $36.26$% under the few-shot learning setting.
@@
For the first challenge, we depict feature distributions and visualize the representations to investigate the underlying cause of why existing methods perform inconsistently for various attacks, including unknown ones and known ones. Figure 2 (b) and (c) depict statistical distributions of the two unknown attacks for the SVM-based model during testing. We observe that feature distributions of MITM attacks are entangled, while the ones of DDoS are more separated. It can be inferred that statistical distributions of traffic features are one of the main underlying causes of performance variations. Separated distributions benefit the unknown attack identification, while entangled ones are indistinguishable and unable to help the NIDS to make accurate decisions. We refer to such a phenomenon as the entangled distribution of statistical features. To analyze the reason for performance variations of acknowledged attacks during testing, we use Pearson correlation heat map [21] to visualize the representations of MITM and DDoS respectively, where the representations are generated by the encoder of E-GraphSAGE. Figure 2 (e) and Figure 2 (f) demonstrate the two correlation maps. Interestingly, we observe that the coefficients of MITM representations are much larger than those of DDoS. We further compare MITM with other attacks, including Backdoor and Dos, and find those high coefficients in the representation will lead to lower intrusion detection scores. We refer to such a phenomenon as the entangled distribution of representational features, which can be considered as another main cause for the degradation of attack classification.
(a) The t-SNE of DIDS
(b) The t-SNE of BSNet
(c) The correlation map of DIDS
(d) The correlation map of BSNet
Figure 4: The correlation map and the t-SNE visualization of representations generated by DIDS and BSNet on CIC-ToN-IoT dataset.
For the second challenge, we observe that the NIDS approaches still perform poorly when combining the state-of-the-art few-shot learning methods, as shown in Fig. 3 (b). For the meta-learning-based MBase [22], we attribute the poor performance to the limited anomaly types in network traffic. Effective meta-learning approaches typically necessitate vast types of few-shot anomalies, which are more aligned with the field of computer vision and natural language processing[23, 24, 25, 26]. Therefore, we focus on the similarity-based method BSNet [27] and uncover why BSNet outperforms DIDS. We depict the t-SNE visualizations of their respective learned representations. As shown in Fig. 4(a) (b), we observe that the representations of BSNet in different attack types are more separated in the latent space. It can be inferred as one of the main causes of higher F1 scores in BSNet. Following this thread, we further depict their representation correlations via correlation heatmaps. We have found that the representations of BSNet are more entangled than DIDS as shown in Fig. 4(c) (d). Aligning with the disentanglement studies in Challenge $1$, we attribute the main obstacle to the performance improvement of BSNet to the entangled representations.
In light of the discussions, we raise two critical questions:
  • $Q1$: How can an intrusion detection model automatically address the first challenges, i.e., two entangled distributions, to benefit the detection of both unknown and known attacks
  • $Q2$: Can a few-shot intrusion detection method automatically learn the representations as separated in latent space, simultaneously as disentangled among representation elements?
@@
Achieving $Q1$ is challenging. To mitigate the issue of the first entangled distribution, we need to differentiate tens and thousands of features involved in real-time network traffic, without prior knowledge of the statistical distributions. Such a problem is largely under-explored in the field of NIDS. For the second entangled distribution, there are some remotely related methods [28, 29, 30] in other fields, including computer vision and natural language processing. However, these approaches mainly focus on object-level representation learning, and hence they are hardly directly applied to intrusion detection to tackle this challenge.
Second, the difficulties of $Q2$ lie in two aspects: 1. tackling $Q2$ involves two optimization terms, a disentanglement term, and a latent space term. How to optimize them into an end-to-end framework is under-explored in NIDS and few-shot learning. 2. As for latent space optimization, existing similarity-based methods [27, 31, 32] mainly focused on utilizing the original-scale information of representations, while ignoring their multi-scale information to help few-shot traffic separated in latent space.
To address the above two questions, we propose a novel model called Disentangled dynamic Intrusion Detection System with Multi-scale Few-shot Learning (DIDS-MFL). Specifically, DIDS-MFL has two critical components: DIDS and MFL. The former DIDS disentangles the statistical flow features with a non-parametric optimization, aiming to automatically separate entangled distributions for representation learning. We refer to this step as statistical disentanglement. Then DIDS further learns to differentiate the representations by a regularization function, aiming to highlight the salient features for specific attacks with smaller coefficients. We refer to this step as representational disentanglement. DIDS finally introduces a novel graph diffusion module that dynamically fuses the graph topology in evolving traffic. The latter MFL is a flexible plug-and-play module for few-shot detection matching multiple NIDS baselines. We first optimize the multi-scale traffic representations in latent space to distinguish attack-specific information. Then we propose an element-wise disentanglement term to highlight the attack-specific information. Finally, we alternately solve them in an end-to-end framework.
Extensive experiments on five benchmarks show the superiority of our DIDS-MFL. The main contributions are:
@@
  • We propose DIDS-MFL, aiming to mitigate the entangled distributions of flow features for NIDS in various practical scenarios, e.g., known, few-shot, and unknown attacks  [^2] More details on the unknown attack detection ability are available in Appendix 14.1.. To the best of our knowledge, we are the first to quantitatively analyze such an interesting problem and empirically reveal the underlying cause in the intrusion detection field.
@@
  • As for DIDS, we present a double disentanglement scheme for differentiating the general features of various attacks and highlighting the attack-specific features, respectively. We additionally introduce a novel graph diffusion method for dynamic feature aggregation.
  • As for MFL, we propose a fusion-based framework by capturing multi-scale information and disentangling representations. We alternately solve them with rigorous derivation for few-shot intrusion detection. As a plug-and-play module, MFL also shows versatility in multiple baselines and downstream tasks.
@@
  • Extensive quantitative and qualitative experiments on various benchmarks demonstrate the effectiveness of DIDS-MFL and provide some helpful insights for effective NIDS in various practical scenarios.
II RELATED WORK
II-A Network Intrusion Detection System
@@
Existing NIDS can be classified into two groups, i.e., signature-based ones [9, 10, 11] and anomaly-based ones [33, 12, 13, 14]. Signature-based intrusion detection systems (IDS) rely on identifying known attack patterns or signatures. These systems are adept at recognizing established threats, where the attack signatures are predefined, employing rule sets to match against network traffic for indicative malicious activities. However, their dependency on signature databases poses a limitation, as they are typically ineffective against novel or zero-day attacks that do not yet have a defined signature. In contrast, the anomaly-based ones do not rely on prior knowledge of attack signatures. They leverage machine learning or statistical techniques to model normal network behavior and flag deviations as potential threats. Anomaly detection holds the advantage of identifying unknown attacks and adapting to new threat patterns. The anomaly-based ones involve statistical methods [2, 3] and deep learning-based ones[15, 16]. Early deep learning studies model the traffic as independent sequences [34, 35, 36, 37]. Recent popular studies rely on GCN to aggregate traffic information [38, 39, 40]. Most related to our work is Euler [41], which builds a series of static graphs based on traffic flow and then performs information aggregation. Our DIDS differs from the above methods in two aspects: 1) We build dynamic graphs rather than static ones, and such a dynamic aggregation can capture fine-grained traffic features for attack detection. Furthermore, we fuse the network layer information into our graph aggregation. 2) We introduce a double disentanglement scheme, including statistical disentanglement and representational one, to benefit the detection of both known and unknown attacks.
II-B Disentangled Representation Learning
@@
Disentanglement aims to learn representations that separate the underlying explanatory factors responsible for variation in the data. Previous studies [42, 43, 44, 45] focus on the generative models by employing constraints on the loss functions, such as $\beta$-VAE [46] modifying the VAE framework to emphasize the independence of latent variables. Beta-VAE achieves disentanglement by introducing a trade-off term between reconstruction and KL divergence, while FactorVAE[47] introduces a total correlation penalty. CausalVAE[48] takes a significant step by integrating causal reasoning, allowing for interventions in the generative process. Disentangled Graph Neural Networks are another active area, where methods aim to learn independent node representations decoupled from the graph structure, e.g., GD-GAN employs a GAN framework to disentangle the latent factors of variation in graph data, allowing for controllable generation of new graph structures and node features. Some recent approaches [49, 50, 51, 52] capture the intrinsic factors in graph-structured data. Most related to our work is DisenLink  [52], which disentangled the original features into a fixed number of factors, with selective factor-wise message passing for better node representations. While our DIDS uses a double disentanglement method, which first disentangles the statistical features via non-parametric optimization, and then learns to highlight the attack-specific features with a regularization.
II-C Dynamic Graph Convolution Networks

### v3

# 3D-IDS: Doubly Disentangled Dynamic Intrusion Detection

@@ Abstract.
Network-based intrusion detection system (NIDS) monitors network traffic for malicious activities, forming the frontline defense against increasing attacks over information infrastructures. Although promising, our quantitative analysis shows that existing methods perform inconsistently in declaring various unknown attacks (e.g., $9\%$ and $35\%$ F1 respectively for two distinct unknown threats for an SVM-based method) or detecting diverse known attacks (e.g., $31$% F1 for the Backdoor and $93$% F1 for DDoS by a GCN-based state-of-the-art method), and reveals that the underlying cause is entangled distributions of flow features. This motivates us to propose 3D-IDS, a novel method that aims to tackle the above issues through two-step feature disentanglements and a dynamic graph diffusion scheme. Specifically, we first disentangle traffic features by a non-parameterized optimization based on mutual information, automatically differentiating tens and hundreds of complex features of various attacks. Such differentiated features will be fed into a memory model to generate representations, which are further disentangled to highlight the attack-specific features. Finally, we use a novel graph diffusion method that dynamically fuses the network topology for spatial-temporal aggregation in evolving data streams. By doing so, we can effectively identify various attacks in encrypted traffics, including unknown threats and known ones that are not easily detected. Experiments show the superiority of our 3D-IDS. We also demonstrate that our two-step feature disentanglements benefit the explainability of NIDS.
Intrusion Detection; Anomaly Detection; Network Security
journalyear: 2023copyright: acmlicensedconference: Proceedings of the 29th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 6–10, 2023; Long Beach, CA, USA.booktitle: Proceedings of the 29th ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD ’23), August 6–10, 2023, Long Beach, CA, USAprice: 15.00isbn: 979-8-4007-0103-0/23/08doi: 10.1145/3580305.3599238ccs: Security and privacy Mobile and wireless security
1. Introduction
@@
Figure 1. Illustration of two network attacks DoS and MITM. DoS floods the target with massive traffic to overwhelm an online service, and MITM eavesdrops on the communication between two targets and steals private information. An NIDS can be easily deployed in a single location to collect statistical features and alert the administrator for potential threats.
Network attacks (Bhushan and Sahoo, 2018), such as password cracking (Kelley et al., 2012), man-in-the-middle attacks (MITM) (Cekerevac et al., 2017), and denial of service (Dos) (Prasad, 2014), refer to unauthorized attempts on various digital assets in an organization’s networks, to steal data or perform malicious actions. Network attacks can also be broadly regarded as network anomalies (Li et al., 2006), as they may involve different characteristics from most other traffics. It was reported that $31\%$ of companies worldwide are now being attacked at least one time a day (Wueest, 2014), due to the growing trends of mobile online business. This urgently calls for an intelligent system that can help administrators automatically filter out these network anomalies in a huge amount of online traffics. A network-based intrusion detection system (NIDS) (Kumar et al., 2016), which monitors network traffic and identifies malicious activities, facilitates administrators to form the frontline defense against increasing attacks over information infrastructures (e.g., sensors and servers). Hence, NIDS is widely applied in many information systems of governments and e-commercial business sectors (Zhao et al., 2009). Figure 1 illustrates that two network attacks threaten information systems, and a NIDS builds a frontline defense against these threats.
@@
Figure 2. Quantitative analysis on CTC-TON-IOT. (a) Comparisons of detecting various attacks, which are regarded as an unknown type in evaluation. Specifically, we train an SVM model without using the data points of these attacks, and evaluate the instances of these attacks on the test set. (b) and (c) show the feature distributions of two attacks, MITM and DDoS, respectively. (d) Comparisons of detecting various known attacks on the previous state-of-the-art deep learning model E-GraphSAGE, (e) and (f) are correlation maps of representations of the two attacks, where the representations are generated by E-GraphSAGE.
Existing NIDS can be categorized into two types, i.e., signature-based ones (Ioulianou et al., 2018; Erlacher and Dressler, 2018; Masdari and Khezri, 2020) and anomaly-based ones (Samrin and Vasumathi, 2017; Van et al., 2017; Aljawarneh et al., 2018; Eskandari et al., 2020). The former detects network attacks based on pre-defined patterns or known malicious sequences stored in a database, such as the number of bytes in traffic. These patterns in the NIDS are referred to as signatures. The latter anomaly-based NIDS learns to track attacks with machine learning techniques. Early statistical (Cekerevac et al., 2017; Prasad, 2014) approaches, such as Support Vector Machine (SVM), Logistic Regression (LR), and Decision Tree (DT), rely on carefully designed handcrafted features to learn classification boundaries. Recent deep learning-based methods (Ahmad et al., 2021; Bhatia et al., 2021) use millions of neural parameters to mine the knowledge underlying the training samples, and have achieved great success in automatically modeling complex correlations for tens and thousands of features. The state-of-the-art E-GraphSAGE (Lo et al., 2022) employs graph convolution networks (GCNs) to learn the feature representations for better prediction.
@@
Although promising, we observe that existing anomaly-based approaches yield inconsistent results in identifying distinct attacks. For statistical methods, Figure 2 (a) demonstrates that the detection performance of an SVM-based method (Ioannou and Vassiliou, 2021) for an unknown attack can be as low as $9\%$ in terms of F1 on CTC-ToN-IOT (Moustafa, 2021), a popular benchmark for NIDS. While the model achieves $40\%$ F1 in declaring another unknown threat (DDoS) on the same benchmark. Regarding the deep learning-based methods, Figure 2 (d) demonstrates that E-GraphSAGE achieves a lower than $20$% F1 score for MITM attacks, and a higher than $90$% F1 score for DDoS on the CTC-ToN-IOT dataset.
@@
To investigate the underlying cause of why existing methods perform inconsistently for various attacks, including unknown ones and known ones, we depict feature distributions and visualize the representations. Figure 2 (b) and (c) depict statistical distributions of the two unknown attacks for the SVM-based model during testing. We observe that feature distributions of MITM attacks are entangled, while the ones of DDoS are more separated. It can be inferred that statistical distributions of traffic features are one of the main underlying causes of performance variations. Separated distributions benefit the unknown attack identification, while entangled ones are indistinguishable and unable to help the NIDS to make accurate decisions. We refer to such a phenomenon as the entangled distribution of statistical features. To analyze the reason for performance variations of acknowledged attacks during testing, we use Pearson correlation heat map (Cohen et al., 2009) to visualize the representations of MITM and DDoS respectively, where the representations are generated by the encoder of E-GraphSAGE. Figure 2 (e) and Figure 2 (f) demonstrate the two correlation maps. Interestingly, we observe that the coefficients of MITM representations are much larger than those of DDoS. We further compare MITM with other attacks, including Backdoor and Dos, and find those high coefficients in the representation will lead to lower intrusion detection scores. We refer to such a phenomenon as the entangled distribution of representational features, which can be considered as another main cause for the degradation of attack classification.
@@
In light of the above discussion, a critical question arises: ”How can an intrusion detection model automatically address the above issue, i.e., two entangled distributions, to benefit the detection of both unknown and known attacks”. Achieving this goal is challenging. To mitigate the issue of the first entangled distribution, we need to differentiate tens and thousands of features involved in real-time network traffic, without prior knowledge of the statistical distributions. Such a problem is largely under-explored in the field of NIDS. For the second entangled distribution, there are some remotely related methods (Liu and Guo, 2019; Brodbeck and Simon, 2020; Sharma and Mir, 2020; Guo et al., 2019) in other fields, including computer vision and natural language processing. However, these approaches mainly focus on object-level representation learning, and hence they are hardly directly applied to intrusion detection to tackle this challenging problem.
We answer the above question and address the two challenging problems with a novel method called 3D-IDS (Doubly Disentangle Dynamic IDS). Specifically, we first disentangle the statistical flow features with a non-parametric optimization based on mutual information, so that entangled distributions can be automatically separated for modeling the general features of various attacks. We refer to this step as statistical disentanglement. Then we further learn to differentiate the representations by a loss function, aiming to highlight the salient features for specific attacks with smaller coefficients. We refer to this step as representational disentanglement. We finally introduce a novel graph diffusion method that fuses the graph topology for spatial-temporal aggregation in evolving traffic. Extensive experiments on five benchmarks show the superiority of our 3D-IDS. The main contributions of this paper are as follows.
@@
  • We propose 3D-IDS, a novel method that aims to mitigate the entangled distributions of flow features for NIDS. To the best of our knowledge, we are the first to quantitatively analyze such an interesting problem and empirically reveal the underlying cause in the field of intrusion detection.
@@
  • We present a double-feature disentanglement scheme for modeling the general features of various attacks and highlighting the attack-specific features, respectively. We additionally introduce a novel graph diffusion method for better feature aggregation.
@@
  • We conduct extensive quantitative and qualitative experiments on various benchmarks and provide some helpful insights for effective NIDS.
2. RELATED WORK
2.1. Network Intrusion Detection System
@@
Existing NIDS can be classified into two groups, i.e., signature-based ones (Ioulianou et al., 2018; Erlacher and Dressler, 2018; Masdari and Khezri, 2020) and anomaly-based ones (Samrin and Vasumathi, 2017; Van et al., 2017; Aljawarneh et al., 2018; Eskandari et al., 2020). The latter involves statistical methods (Cekerevac et al., 2017; Prasad, 2014) and deep learning-based ones(Ahmad et al., 2021; Bhatia et al., 2021). Early deep learning studies model the traffic as independent sequences (Ma et al., 2020; Hu et al., 2021; Li et al., 2022; Doriguzzi-Corin et al., 2020). Recent popular studies rely on GCN to aggregate traffic information (Kalander et al., 2020a; Cong et al., 2021; Song et al., 2020). Most related to our work is Euler (King and Huang, 2022), which builds a series of static graphs based on traffic flow and then performs information aggregation. Our 3D-IDS differs from the above methods in two aspects: 1) We build dynamic graphs rather than static ones, and such a dynamic aggregation can capture fine-grained traffic features for attack detection. Furthermore, we fuse the network layer information into our graph aggregation. 2) We also introduce a two-step disentanglement scheme, including statistical disentanglement and representational one, to benefit the detection of both known and unknown attacks.
2.2. Disentangled Representation Learning
@@
Disentanglement aims to learn representations that separate the underlying explanatory factors responsible for variation in the data. Previous studies (Bengio et al., 2013; Chen and Batmanghelich, 2020; Gilpin et al., 2018; Gondal et al., 2020; Locatello et al., 2020; Lipton, 2018; Creager et al., 2019) focus on the generative models by employing constraints on the loss functions, such as $\beta$-VAE (Higgins et al., 2017). Some recent approaches (Suter et al., 2019; Ridgeway and Mozer, 2018; Träuble et al., 2021; Zhou et al., 2022) have studied disentangled GNNs to capture the intrinsic factors in graph-structured data. Most related to our work is DisenLink  (Zhou et al., 2022), which disentangled the original features into a fixed number of factors, with selective factor-wise message passing for better node representations. While our 3D-IDS uses a double disentanglement method, which first disentangles the statistical features via non-parametric optimization, and then learns to highlight the attack-specific features with a regularization.
2.3. Dynamic Graph Convolution Networks